Perform an engineering economy analysis of multiple alternatives –decide between renting or buying a house. Determine the present worth of both investments. Given Information: Buy Rent Cost $500,000 Initial Payment Down Payment: 10% of cost Monthly Payments Security Deposit: $3,500
Answer #1

We will calculate the NPV of cash outflow of both the options and select the option which have least cost:

Option 1: Buying the property

Cost= 500000

Down payment= 10% of Cost i.e $50000

Remaining amount i.e. $ 450000 will be paid as EMI for next 15 years @ $% interest rate compounding monthly:

Calculation of EMI:

[P x R x (1+R)^N] / [(1+R)^N-1]

Here,

P= Loan Amount =450000

R= Monthly Interest rate = 0.04/12 = 0.0033

N= Total No. of installments = 15 x 12 = 180 Months

EMI= [450000 × 0.0033 x (1 + 0.0033)^180 ] / (1 + 0.0033)^180-1

EMI = 1485 × (1.0033)^180 / 1.8094-1

EMI =1485 × 1.8094 / 0.8094

EMI = 2687 / 0.8094

Equal Monthly Installment = $ 3320

NPV of outflow of option 1:

NPV of Down Payment: 50000 x 1 = 50000

NPV of EMI (Compounding factor calculated @ 6%)= 3320 x 118.5035 = 393432

NPV of Insurance =

year (A) Insurance Amount (B) Discount Factor(C ) NPV (D)= B X C
0 400 1 400.00
1 406 0.9434 383.02
2 412.09 0.8900 366.76
3 418.27135 0.8396 351.19
4 424.54542 0.7921 336.28
5 430.913602 0.7473 322.00
6 437.377306 0.7050 308.33
7 443.937965 0.6651 295.24
8 450.597035 0.6274 282.71
9 457.35599 0.5919 270.71
10 464.21633 0.5584 259.22
11 471.179575 0.5268 248.21
12 478.247269 0.4970 237.67
13 485.420978 0.4688 227.58
14 492.702292 0.4423 217.92
Total NPV 4506.85

Total NPV of option 1 = 50000+393432+4507 = $ 447939

option 2: Renting

NPV of outflow under option 2: $ 357559.25 (calculation made in excel spreadsheet):

NPV of Security Deposit= 3500 x 1 = 3500

NPV of rent:

Year Monthly Outflow(B) Sum of Discount Factor(C ) Sum of NPV (D)= B X C
1 2500.00 11.62 29047.33
2 2575.00 10.94 28180.63
3 2652.25 10.31 27339.79
4 2731.82 9.71 26524.04
5 2813.77 9.15 25732.63
6 2898.19 8.61 24964.83
7 2985.13 8.11 24219.94
8 3074.68 7.64 23497.28
9 3166.93 7.20 22796.18
10 3261.93 6.78 22116.00
11 3359.79 6.39 21456.11
12 3460.58 6.02 20815.92
13 3564.40 5.67 20194.82
14 3671.33 5.34 19592.26
15 3781.47 5.03 19007.67
Grand Total 118.50 355485.44
NPV of Security Deposit refunded back = 3500 x 0.4075 =1426.19


Total outflow NPV of option 2 = 3500+355485.44-1426.19 =357559.25

As the cash outflow NPV of option 2 i.e renting is less as compare to option 1 i.e. Buying the property, therefore, option 2 will be preferred.